OpenStudy (anonymous):

having issues....thought i understood it, but i keep getting the wrong answer. can someone explain what I'm doing wrong here? parametric curve x=5t+lnt y=6t-lnt find d^2y/dx^2

OpenStudy (anonymous):

so far i have dy/dx = 6t-1/5t+1

OpenStudy (anonymous):

d^2y/dx^2=\[\frac{ d/dt(\frac{ 6t-1 }{ 5t+1 } }{ 5t+1 }\]

OpenStudy (anonymous):

d/dt = 30t+6-30t+5/(5t+1)^2=11/(5t+1)^2 all over another 5t+1 giving 11/(5t+1)^3

OpenStudy (anonymous):

but webwork says it's wrong...what did i do?

OpenStudy (rational):

\(\large \frac{d^2y}{dx^2} = \frac{d/dt (dy/dx)}{dx/dt}\)

hartnn (hartnn):

denominator = 5 + 1/t and not just 5t+1 right ?

OpenStudy (rational):

\(\large \frac{d^2y}{dx^2} = \frac{d/dt (\frac{6t-1}{5t+1})}{5 + 1/t}\)

hartnn (hartnn):

because dx/dt is 5+1/t

OpenStudy (anonymous):

http://www.youtube.com/watch?v=7ANhq51wjWM

OpenStudy (anonymous):

so that video explaining it is wrong?

OpenStudy (rational):

first slide in that video looks correct oly ? it uses the same formula we're using eh ?

OpenStudy (anonymous):

it doesn't have the extra t...i did the steps the same way and there is no extra t...it just divides everything by dx which is 5t+1. you guys are saying divide by 5t+1/t

hartnn (hartnn):

what did you get as dx/dt =... ?

OpenStudy (rational):

\(\large \frac{dx}{dt} = 5 + \frac{1}{t}\)

OpenStudy (rational):

once u have dy/dx, to take the second derivative, you need to do it in parametric way.. find dy'/dt, and divide it wid dx/dt ** formula above

OpenStudy (anonymous):

the derivative of x is 5+1/t or 5t+1. all these dy, dx, dt, confuses the tar out of me. d/dx is derivative of x, right?

OpenStudy (rational):

**the derivative of x is 5+1/t or 5t+1 who ate the bottom t ? :)

OpenStudy (rational):

5 + 1/t = (5t+1)/t

hartnn (hartnn):

5+1/t = (5t+1)/t

OpenStudy (anonymous):

no one...it's just 5+1/t...times the whole thing by t to make it simpler to deal with gets you 5t+1

OpenStudy (anonymous):

5+(1/t) *t is 5t+1

OpenStudy (rational):

ugh by that logic, 1+1/3 is same as 3*1 + 1 is it

hartnn (hartnn):

why ...*t ?

OpenStudy (anonymous):

to get rid of the fraction...(5+1/t) *t distribute the t = 5t+t/t or 5t+1

OpenStudy (rational):

to simplify and get rid of fraction, u need to multiply wid 1 = t/t both top and bottom u need to multiply wid t ok

OpenStudy (rational):

\(\large 5 + \frac{1}{t}\) \(\large 5\frac{t}{t} + \frac{1}{t}\) \(\large \frac{5t+1}{t}\)

OpenStudy (rational):

sorry, here you cant get rid of fraction as such...

OpenStudy (rational):

looks u have applied the formula correctly, just made an error during simplifying fracitons :o

OpenStudy (anonymous):

i got that rewrite directly from a derivative calculator....

OpenStudy (anonymous):

it said 6-1/t / 5+1/t was the same as 6t-1/5t+1.

OpenStudy (rational):

its same, but in ur derivative calculator page, you're just finding \(\large \frac{d}{dt} (\frac{dy}{dx})\)

OpenStudy (rational):

you need to divide watever u get above by dx/dt ok

OpenStudy (anonymous):

i'm so confused....so it's the derivative of the dy/dx times what over what? i don't understand what d/dt(dy/dx)/dx/dt...that's just too many d's t's x's and y's...

OpenStudy (rational):

lets start from scratch. it wont take more than 10 minutes to digest this for u

OpenStudy (anonymous):

yeah, i understand "derivative" and f prime and all that, but these notations confuse me. thank you for being patient with me, @rational

OpenStudy (rational):

\(x=5t+\ln t\) \(y=6t-\ln t \) \(\large \frac{dx}{dt} = 5 + \frac{1}{t}\) \(\large \frac{dy}{dt} = 6 - \frac{1}{t}\) \(\huge \frac{dy}{dx} = \frac{\frac{dy}{dt}}{\frac{dx}{dt}}\) \(\huge ~~~= \frac{6 - \frac{1}{t}}{5 + \frac{1}{t}}\) \(\huge ~~~= \frac{6t-1}{5t+1}\)

OpenStudy (rational):

this far, you have worked already.

OpenStudy (rational):

it takes time to get use to, dont wry... two more problems like these... and u wil feel confy :)

OpenStudy (rational):

next lets find \(\large \frac{d^2y}{dx^2}\)

OpenStudy (rational):

\(\huge \frac{d^2y}{dx^2} = \frac{\frac{d}{dt} (\frac{dy}{dx})}{\frac{dx}{dt}}\)

OpenStudy (anonymous):

ok, with you so far.

OpenStudy (rational):

plug things in above formula

OpenStudy (rational):

\(\huge \frac{d^2y}{dx^2} = \frac{\frac{d}{dt} (\frac{dy}{dx})}{\frac{dx}{dt}}\) \(\huge \frac{d^2y}{dx^2} = \frac{\frac{d}{dt} (\frac{6t-1}{5t+1})}{5+\frac{1}{t}}\)

OpenStudy (rational):

find the derivative on numerator and simplify.

OpenStudy (rational):

u knw (f/g)' formula right

OpenStudy (anonymous):

ok, so (11/(5t+1)^2) / (5+1/t)

OpenStudy (rational):

thats correct ! you may simplify more if u want

OpenStudy (rational):

\(\huge \frac{d^2y}{dx^2} = \frac{\frac{d}{dt} (\frac{dy}{dx})}{\frac{dx}{dt}}\) \(\huge \frac{d^2y}{dx^2} = \frac{\frac{d}{dt} (\frac{6t-1}{5t+1})}{5+\frac{1}{t}}\) \(\huge \frac{d^2y}{dx^2} = \frac{ \frac{11}{(5t+1)^2}}{5+\frac{1}{t}}\) \(\huge \frac{d^2y}{dx^2} = \frac{ \frac{11}{(5t+1)^2}}{\frac{5t+1}{t}}\) \(\huge \frac{d^2y}{dx^2} = \frac{11t}{(5t+1)^3}\)

OpenStudy (anonymous):

thanks. i got the right answer. finally. test friday.

OpenStudy (rational):

:) good luck !